OpenStudy (anonymous):

Three squares are arranged so as to form a triangle in between them. The area of square X is 16 square inches, the area of square Y is 9 inches. What is the length of a side of square z? Anyone?

OpenStudy (anonymous):

Ok, so I'm confused. Can you walk me through it?

Directrix (directrix):

Look at the diagram The square with area 16 has side 4. The square with area 9 has side 3. 4 ^2 + 3^2 = c^2 16 + 9 = c^2 25 = c^2 5 = c